Q8. The pressure of the gas when measuring its volume as described in the lab is approximately 1 atm, and the temperature is approximately 300 K. Using the volume of gas you measured in the lab, how many moles of CO2 did you observe as reaction products: (volume of gas measured was 400 mL for 1/2tsp and 300 mL for 1/4tsp.

  1. From ¼ tsp baking soda?

_____________________________________ moles

  1. From ½ tsp baking soda?

_____________________________________ moles

Q9. Calculate your percent error for each:

_____________________________________ g/mol

  1. From ¼ tsp baking soda?

_____________________________________ %

  1. From ½ tsp baking soda?

_____________________________________ %
Answer #1

(a) for (1/4)tsp

PV = nRT

1atm × 300×10-3L = n × 0.0821atm-L/K.mol × 300K

n = 300×10-3 /(0.0821×300) = 0.01218 moles

(b)for (1/2) tsp

PV=nRT

1atm × 400×10-3L = n × 0.0821atm-L/K.mol × 300K

n = 0.01624moles

Q9:

(c) mass of 1/4 tsp = 5/4 g = 1.25g

Molar mass of baking soda(NaHCO3) = 84g/mol

Moles of NaHCO3 = 1.25g/ 84g/mol = 0.01488moles

Moles of NaHCO3 is equal to the moles of the CO2 gas released.

Moles of CO2 = 0.01488moles

% error =[ (0.01488 - 0.01218)/0.01488]×100 = 18.14%

(d) moles of NaHCO3 in (1/2)tsp = 2× moles of NaHCO3 in (1/4)tsp = 0.01488×2 = 0.2976moles

%error = [(0.02976 - 0.01624)/0.02976]×100 = 45.43%